# Towards Robust Surrogate Models: Benchmarking Machine Learning Approaches to Expediting Phase Field Simulations of Brittle Fracture

### v1

@@ 2.2 Dataset specifications
The version of our dataset published in conjunction with this paper contains 6,000 two dimensional quasi static simulations of phase field modeling for brittle fracture. We simulate domains with internal initial cracks under biaxial tension and shear boundary conditions. For the shear case, the top and bottom boundaries are also subject to $\phi=0$ Dirichlet boundary condition as in many examples without this condition, the fracture initiated and propagated along these boundaries, resulting in uninformative fracture patterns that are not useful for training ML models. We run all three energy decomposition methods (1000 initial conditions $\times$ 2 boundary conditions $\times$ 3 energy decomposition methods $=$ 6000 simulations in total). Schematics of representative geometries and applied boundary condition are shown in Fig. 5a. The overall dataset structure is as follows. First, each simulation is categorized under one of 6 sub-categories named as case-energy_decomp for example: tension-spect. Each sample is saved under the random seed that was used for generating the random internal cracks in order to enhance reproducibility. The random parameters in each sample are the number of internal cracks which is an integer between 10-20, and the coordinates and orientations of the crack inside the domain. All of the cracks have the same length $l=0.25$ mm. For each sample, at 100 simulation steps, we save the phase field, the x and y components of the displacement field on the mesh, the force displacement data on the support edges, and the data is saved in HDF5 format. The simulation parameters including Young’s modulus $E$, Poisson’s ratio $\nu$, critical energy release rate $G_{c}$, characteristic length $l_{0}$ and the mesh resolution are listed in Table 1. A uniform structured mesh with bi-linear elements were used with 4 quadrature points in each element.
@@
$l_{c}$ (Length scale) 0.01 mm
@@
The dataset can be accessed in both a $128\times 128$ grid format, and an original mesh resolution format. The $128\times 128$ format offers a good balance of information and data size maintaining a standard resolution for ML applications and ease of use, while the original mesh resolution format enables flexibility. In Fig. 6, we also show that different energy decompositions sometimes – but not always – result in significantly different outcomes.
@@ 2.3.1 Physics Informed Neural Networks
The goal of training a Physics-Informed Neural Network (PINN) is to approximate the solution to a partial differential equation (PDE). The concept was introduced in [7], and since then there has been an explosion of related work in the literature [83, 84, 85, 86, 87]. In some cases, neural networks have been used to minimize the residual of the strong form of the PDE, as in [7], while in others, the variational form of the PDE has been used, as in [85, 88, 79]. There have been multiple recent attempts to apply PINNs to phase-field models of brittle fracture, for example [46, 47, 89, 40, 39, 48, 90]. In some of these studies, to mitigate the computational cost of training a PINN for every loading step, different transfer learning approaches have been explored. For instance, the trained weights from one loading step can be used as a good initialization point for the next step [40], or the weights of all layers except the last can be kept fixed after training during the first step, and then only the weights of the last layer of the feedforward network are retrained [39]. A typical PINN architecture consists of multilayer perceptrons (MLPs), and usually weighted combinations of multiple loss terms that are mainly physics-informed as a form of a PDE residual and can also have supervised data-driven terms, on a set of collocation points inside the domain or on the boundary [7]. PINN based models are usually trained for a great number of iterations to drive the residual toward zero.
@@
For our representative baseline model, we explore the process of minimizing the PINN loss function for only a single set of boundary conditions and initial cracks over a predefined number of loading steps. In our preliminary explorations (not reported in detail in this paper), we attempted both the vanilla PINN formulation using the DeepXDE Python package [91], and the DRM model developed by Manav et.al.  [40]. The DRM uses the quadrature points of the mesh as the collocation points, and is formulated to minimize the energy functional of the variational problem [46, 40]. For the main PINN baseline model presented in this paper, we have used the PINN architecture and code provided by Manav et al. [40] as a representative example from the recent literature of a PINN formulation with phase field fracture as the target application. We chose this approach because the implementation was well documented and openly available and was benchmarked on the crack coalescence case which aligned closely with our dataset. To make a fair comparison, a uniform unstructured mesh with the same mesh size as the FEA solution described in Section 2.2 with triangular elements was provided as the input to the model. After experimenting with different hyperparameters such as width and depth of the hidden layers, we settled on 6 hidden layers, each having 100 neurons and with characteristic length set to 0.01 mainly because other configurations were computationally prohibitive. As stated in the original paper, training a model can take up to an order of magnitude longer for much simpler crack pattern cases. We note briefly that though this method has demonstrated efficacy on simpler initial crack patterns (e.g., single-edge notched specimen under tension and shear and crack coalescence), it has not been rigorously tested on more complex patters such as the ones provided in our dataset.
@@ 2.3.2 Fourier Neural Operator (FNO)
In this work, we have followed the PDEBench [35] implementation and used the same architecture and hyperparameters as the original paper by Li et al. [92]. Specifically, this implementation has four Fourier layers, with the modes in each direction set to 12 with Gaussian Error Linear Unit (GELU) activation functions. We trained the model for 300 epochs with the mean squared error (MSE) loss and the Adam optimizer [96] setting the initial learning rate to 0.001 with a scheduler set to halve the learning rate every 100 epochs, and a batch size of 4. The FNO model was trained autoregressively using three channels for the input (phase field, displacement field in x direction, and displacement field in y direction) with ten steps for autoregressive rollout for five different seeds. In addition, we 0-1 normalized the input and output of the displacement fields of the FNO for each loading step based on the applied displacement on the boundaries.
@@ 2.3.3 UNet
Here we use the UNet architecture implemented in [82]. Due to their ease of implementation and popularity in the broader literature [35, 34], UNets are a natural choice as a baseline model [97]. After preliminary exploration of multiple potential UNet architectures and loss functions, we settled on the UNet architecture used in previous work by our group for investigating model calibration for classification [82]. In contrast to our FNO implementation, we did not train the UNets autoregressively because in our initial attempts, autoregressive training led to unstable behavior. And, following preliminary exploration, we only trained the UNets with the phase field channel. Because the UNet struggled with the limited number of training samples, we applied data augmentation such as random rotations and horizontal and vertical flips on the fly to improve its performance. In contrast, the FNOs did not require such augmentation, since they were trained autoregressively over the time steps and did not suffer from the same limitation. We set the batch size to 16 for 100 epochs using the Adam optimizer with the learning rate set to 0.0001. For the UNet we used the sum of Dice and Focal loss with $\gamma=2$ and $\alpha=0.8$ [98] with equal weights and we used 0.4 threshold for binarizing the ground truth. The same dataset and data split was used for the UNet as the FNO.
@@ 2.4 Ensembling methods
We trained both the FNO and UNet models with five different random seeds. To showcase performance in this paper, we used the dataset with bi-axial tension boundary condition and the spectral decomposition energy split with a $128\times 128$ input resolution. The dataset was split into $80\%$ training data, with the remaining $20\%$ split into $60\%$ as the meta train split to train the stacking ensemble, $20\%$ as the meta validation set for tuning the hyperparameters of the stacking model, and the final $20\%$ used as the meta test set. As seen in Fig. 10, models trained with different random seeds often predict different results. Compellingly, and consistent with our previous work in the area [82], the multiple different results often look “plausible” which motivates the general idea of exploring methods that can combine these different results with an eye towards future exploration from an uncertainty quantification perspective [99]. In line with the goals of this paper, we provide baseline results from performing straightforward ensemble averaging across the five model seeds using majority and unweighted averaging also known as hard and soft voting respectively [100, 101] and stacking methods [102]. A brief explanation of each approach follows.
@@
Hard voting is a well known ensembling method also referred to as “majority voting.” In this simple approach, we aggregate the class label predictions from different models, and the majority vote determines the final prediction. Soft voting is another well known and straightforward ensembling method that averages the class label probabilities across different models [100]. It has been observed that even using these simple ensembling methods can improve the performance of the models such as UNet [82], and different random initializations of the models can lead to learning different patterns and modes of the data and ensembling the results of them would reduce the variance of the final prediction [103, 5, 101]. Finally, stacking, originally introduced in [102], is an ensembling method that trains a meta-learner on the predictions of the base models to learn a linear or nonlinear combination of their outputs, producing a final prediction for the class label. To perform our baseline model of stacking, we used a simple architecture with 81 parameters that is comprised of two 2D convolutional layer with kernel size 1 and 3 to denoise and capture local features on all five channels and then a ReLU activation function and a final 2D convolutional layer that aggregates all five channels into the final output. We have observed that in this small network the initialization is very important to get a good result. As the probability of having a cracked pixel inside the domain is very low (approximately $2\%$) we initialized the weights of the first and the last layer using Xavier initialization [29]. Additionally we initialize the bias of the first layer to 0.2, enabling the model to start from a soft voting state in which each of the five sub models is given equal weights. The $3\times 3$ depthwise layer is initialized with the identity stencil with no bias. The reasoning behind this initialization was so that the initial pass through the network would start from a stable point and to prevent dead ReLU activations. We trained the stacker model for 50 epochs on the meta train split of both UNet and FNO predictions. As the stacking method in its most basic form can approximate soft voting by learning equal weights for each channel, it should be noted that its true potential lies in extending it with added contextual windows and nonlinearities. This allows the network to learn a nonlinear combination of the sub models’ outputs and to learn from weaknesses and strengths of them across different regions of the field. Though many more advanced ensembling techniques exist [104, 105, 106] and are considered “state of the art,” we consider them beyond the scope of this paper where our goal is to provide a set of baseline models.
@@
Figure 7: Figure showing: a) The initial condition and the resulting FEA simulation under bi-axial tensile loading and b) The results of training the PINN model (12 separate models) with three energy decomposition methods across four different seeds. We note that the individual PINN models are converging to different crack patterns. Though some of these models show plausible results, none were able to predict the correct crack pattern.
@@ 3 Results and Discussion
Figure 8: This figure presents the predicted phase field and displacement fields for each FNO model that was trained on the dataset. It shows the results for a representative sample at the $50^{th}$ percentile of performance with a Dice score of 0.6869. The top row shows the ground truth and each subsequent row shows the results of each FNO seed. Additional examples are shown in Fig. 13
@@ 3.2 Fourier Neural Operator Baseline
In Fig. 9a, we plot a histogram of the Dice score across all 40 test samples evaluated with the trained FNO models. As indicated by the vertical line labeled “Dice Mean Seed 1-5” in the Fig. 9a histogram, the average Dice score across the test set and across all seeds is 0.6208. We can see in Fig. 8 that FNO models predicted a visually less well-defined and smoother phase field solution when compared to the FEA Ground Truth. A plausible explanation for this outcome is related to a well-known limitation of neural networks referred to as spectral bias, which indicates the preference of neural networks to learn low frequency features at the expense of high frequency features [23, 107, 108, 24]. Another interesting observation from Fig. 8 is that although the predicted displacement fields may deviate from the ground truth, they remain consistent with their corresponding phase fields. This suggests that the models have captured some underlying physical relationships through training on this dataset.
Looking ahead, there are several promising directions to improve upon these baseline models. For example exploration of various data augmentation strategies, ranging from traditional techniques to more advanced generative AI based approaches. application of different data augmentation techniques from simpler methods to generative AI based methods. As shown in Fig. 9b, which presents crack pattern predictions ranked from best to worst based on the soft voting method (discussed further in Section 2.4), we observe that examples with lower Dice scores typically have large quantitative and qualitative deviations from the ground truth. Thus, although the top-performing predictions (e.g., those in the upper row of Fig. 9b) show strong agreement with the FEA ground truth, the substantial discrepancies in lower-performing cases highlight significant room for improvement.

### v2

@@ 2.2 Dataset specifications
The version of our dataset published in conjunction with this paper contains 6,000 two dimensional quasi static simulations of phase field modeling for brittle fracture. We simulate domains with internal initial cracks under the Bi-axial tension and Shear boundary conditions. For the Shear case, the top and bottom boundaries are also subject to $\phi=0$ Dirichlet boundary condition as in many examples without this condition, the fracture initiated and propagated along these boundaries, resulting in uninformative fracture patterns that are not useful for training ML models. We run all three energy decomposition methods (1000 initial conditions $\times$ 2 boundary conditions $\times$ 3 energy decomposition methods == 6000 simulations in total). Schematics of representative geometries and applied boundary condition are shown in Fig. 5a. The overall dataset structure is as follows. First, each simulation is categorized under one of 6 sub-categories named as case-energy_decomp for example: tension-spect. Each sample is saved under the random seed that was used for generating the random internal cracks in order to enhance reproducibility. The random parameters in each sample are the number of internal cracks which is an integer between 10-20, and the coordinates and orientations of the crack inside the domain. All cracks have the same length $l=0.25$ mm. As a brief note, this consistent initial crack length was chosen after preliminary investigations in order ensure reliable convergence of the simulations, and avoid situations where crack propagation is dominated by a single large initial crack.
For each sample across all datasets, we saved 100 snapshots of the crack evolution. All simulations used loading increments of $\Delta u=1\times 10^{-6}$ mm. The datasets with the Bi-axial tension boundary condition were simulated for 5,000 steps with a saving frequency of 50. The simulations based on the Shear boundary condition were simulated for 10,000 steps with a saving frequency of 100. In total, for all simulated samples, we saved 100 loading steps. For each saved step, we recorded the phase field, and the $x$ and $y$ components of the displacement field. The force displacement data on the support edges were also recorded for all loading steps. The data is saved in HDF5 format. The simulation parameters including Young’s modulus $E$, Poisson’s ratio $\nu$, critical energy release rate $G_{c}$, characteristic length $l_{0}$ and the mesh resolution are listed in Table 1. A uniform structured mesh with bi-linear elements were used with 4 quadrature points in each element.
@@
$l_{0}$ (Length scale) 0.01 mm
$\Delta u$ (Loading increment) $1\times 10^{-6}$ $m$
@@
Table 2: Simulation specifications for the two boundary conditions and energy decomposition methods. Here, $N_{s}$ denotes the number of samples in each dataset and $N_{t}$ is the number of saved snapshots in each trajectory.
Boundary Condition Decomposition Channels Resolution $N_{s}$ $N_{t}$ Avg. Time (h/sim) Hardware Dataset Size
Bi-axial Tension Spectral $\phi,u,v$ $128\times 128$ 1000 100 2.88 16-core CPU 32.4 GB
Vol.-Dev. $\phi,u,v$ $128\times 128$ 1000 100 2.76 16-core CPU 32.4 GB
Star-convex $\phi,u,v$ $128\times 128$ 1000 100 2.74 16-core CPU 32.4 GB
Shear Spectral $\phi,u,v$ $128\times 128$ 1000 100 5.42 16-core CPU 33.7 GB
Vol.-Dev. $\phi,u,v$ $128\times 128$ 1000 100 5.26 16-core CPU 33.4 GB
Star-convex $\phi,u,v$ $128\times 128$ 1000 100 5.28 16-core CPU 33.6 GB
The dataset can be accessed in both a $128\times 128$ grid format, and an original mesh resolution format. The $128\times 128$ format offers a good balance of information and data size maintaining a standard resolution for ML applications and ease of use, while the original mesh resolution format enables flexibility. In Fig. 6, we also show that different energy decompositions sometimes – but not always – result in significantly different outcomes. All the simulations were run in parallel using 16 cores. The parameters regarding the simulation time, hardware used, and dataset size are listed in Table 2
@@ 2.3.1 Physics Informed Neural Networks
The goal of training a Physics-Informed Neural Network (PINN) is to approximate the solution to a partial differential equation (PDE). The concept was introduced in [10], and since then there has been an explosion of related work in the literature [93, 94, 95, 96, 97]. In some cases, neural networks have been used to minimize the residual of the strong form of the PDE, as in [10], while in others, the variational form of the PDE has been used, as in [95, 98, 89]. There have been multiple recent attempts to apply PINNs to phase field models of brittle fracture, for example [51, 52, 99, 45, 44, 53, 100]. In some of these studies, to mitigate the computational cost of training a PINN for every loading step, different transfer learning approaches have been explored. For instance, the trained weights from one loading step can be used as a good initialization point for the next step [45], or the weights of all layers except the last can be kept fixed after training during the first step, and then only the weights of the last layer of the feedforward network are retrained [44]. A typical PINN architecture consists of multilayer perceptrons (MLPs), and usually weighted combinations of multiple loss terms that are mainly physics-informed as a form of a PDE residual and can also have supervised data-driven terms, on a set of collocation points inside the domain or on the boundary [10]. PINN based models are usually trained for a large number of iterations to drive the residual toward zero. Here we should note that a PINN based model is trained to approximate the solution to a PDE for a specific set of input functions. Operator learning approaches, such as DeepONets [16] and Fourier Neural Operators [101], can utilize physics informed loss terms similar to the PINNs to be able to generalize over a wider range of input functions without the need for retraining.
@@
For our representative baseline model, we explore the process of minimizing the PINN loss function for only a single set of boundary conditions and initial cracks over a predefined number of loading steps. In our preliminary explorations (not reported in detail in this paper), we attempted both the vanilla PINN formulation using the DeepXDE Python package [102], and the DRM model developed by Manav et al.  [45]. The DRM uses the quadrature points of the mesh as the collocation points, and is formulated to minimize the energy functional of the variational problem [51, 45]. For the main PINN baseline model presented in this paper, we have used the PINN architecture and code provided by Manav et al. [45] as a representative example from the recent literature of a PINN formulation with phase field fracture as the target application. We chose this approach because the implementation was well documented and openly available and was benchmarked on the crack coalescence case which aligned closely with our dataset. To make a fair comparison, a uniform unstructured mesh with the same mesh size as the FEA solution described in Section 2.2 with triangular elements was provided as the input to the model. After experimenting with different hyperparameters such as width and depth of the hidden layers, we settled on 6 hidden layers, each having 100 neurons and with characteristic length set to 0.01 in the non-dimensionalized energy functional as described in [45]. As a brief note, training a model with more parameters was not realistic with the computational resources that we have access to. This is consistent with the paper on which this model is based [45], where the authors noted that training a model can take up to an order of magnitude longer for much simpler crack pattern cases.
The trained models use adaptive ReLU as the activation function. Instead of using auto-differentiation, the gradients of the nodal values in each element, approximated by the model, were calculated using linear shape functions. These gradients were then used to approximate the integral within each element via the Gaussian quadrature rule. The L-BFGS [103] optimizer was used for the first step, and the remaining steps used RPROP [104] as the optimizer. We note briefly that though this method has demonstrated efficacy on simpler initial crack patterns (e.g., single-edge notched specimen under tension and shear and crack coalescence), it has not been rigorously tested on more complex patters such as the ones provided in our dataset.
@@ 2.3.2 Fourier Neural Operator (FNO)
In this work, we have followed the PDEBench [38] implementation of basic unrolled autoregressive training and used the same architecture and hyperparameters as the original paper by Li et al. [101]. Specifically, this implementation has four Fourier layers, with the modes in each direction set to 12 with Gaussian Error Linear Unit (GELU) activation functions. We trained the model for 300 epochs with the mean squared error (MSE) loss and the Adam optimizer [108] setting the initial learning rate to 0.001 with a scheduler set to halve the learning rate every 100 epochs, and a batch size of 4. We trained the FNO model using unrolled autoregressive training approach with a 10 step history length. The model is trained for the complete trajectory on the training set. We feed all three channels for the input (phase field, displacement field in x direction, and displacement field in y direction) for five different seeds. In addition, for each loading step, we scaled the input and output displacement fields of the FNO to the range [0, 1], using the applied boundary displacement at that step as the reference maximum value.
@@ 2.3.3 UNet
Here we use the UNet architecture implemented in [92]. Due to their ease of implementation and popularity in the broader literature [38, 37], UNets are a natural choice as a baseline model [109]. After preliminary exploration of multiple potential UNet architectures and loss functions, we settled on the UNet architecture used in previous work by our group for investigating model calibration for classification [92]. In contrast to our FNO implementation, we did not train the UNets autoregressively because in our initial attempts, unrolled autoregressive training led to unstable behavior. And, following preliminary exploration, we only trained the UNets with the phase field channel. Because the UNet struggled with the limited number of training samples, we applied data augmentation such as random rotations and horizontal and vertical flips on the fly to improve its performance. In contrast, the FNOs did not require such augmentation, since they were trained autoregressively over the time steps and did not suffer from the same limitation. We set the batch size to 16 for 100 epochs using the Adam optimizer with the learning rate set to 0.0001. For the UNet we used the sum of Dice and Focal loss with $\gamma=2$ and $\alpha=0.8$ [110] with equal weights and we used 0.4 threshold for binarizing the ground truth. The same dataset and data split was used for the UNet as the FNO.
@@ 2.4 Ensembling methods
We trained both the FNO and UNet models with five different random seeds. To showcase performance in this paper, we used the dataset with the Bi-axial tension boundary condition and the spectral decomposition energy split with a $128\times 128$ input resolution. The dataset was split into $80\%$ training data, with the remaining $20\%$ split into $60\%$ as the meta train split to train the stacking ensemble, $20\%$ as the meta validation set for tuning the hyperparameters of the stacking model, and the final $20\%$ used as the meta test set. As seen in Fig. 11, models trained with different random seeds often predict different results. Compellingly, and consistent with our previous work in the area [92], the multiple different results often look “plausible” which motivates the general idea of exploring methods that can combine these different results in a manner that will enable future study from an uncertainty quantification perspective [111]. In line with the goals of this paper, we provide baseline results from performing straightforward ensemble averaging across the five model seeds using majority and unweighted averaging also known as hard and soft voting respectively [112, 113] and stacking methods [114]. A brief explanation of each approach follows.
@@
Hard voting is a well known ensembling method also referred to as “majority voting”. In this simple approach, we aggregate the class label predictions from different models, and the majority vote determines the final prediction. Soft voting is another well known and straightforward ensembling method that averages the class label probabilities across different models [112]. It has been observed that even using these simple ensembling methods can improve the performance of the models such as UNet [92], and different random initializations of the models can lead to learning different patterns and modes of the data and ensembling the results of them would reduce the variance of the final prediction [57, 8, 113]. Finally, stacking, originally introduced in [114], is an ensembling method that trains a meta-learner on the predictions of the base models to learn a linear or nonlinear combination of their outputs, producing a final prediction for the class label. To perform our baseline model of stacking, we used a simple architecture with 81 parameters that is comprised of two 2D convolutional layer with kernel size 1 and 3 to denoise and capture local features on all five channels and then a ReLU activation function and a final 2D convolutional layer that aggregates all five channels into the final output. We have observed that in this small network the initialization is very important to get a good result. As the probability of having a cracked pixel inside the domain is very low (approximately $2\%$) we initialized the weights of the first and the last layer using Xavier initialization [32]. Additionally we initialize the bias of the first layer to 0.2, enabling the model to start from a soft voting state in which each of the five sub models is given equal weights. The $3\times 3$ depthwise layer is initialized with the identity stencil with no bias. The reasoning behind this initialization was so that the initial pass through the network would start from a stable point and to prevent dead ReLU activations. We trained the stacker model for 50 epochs on the meta train split of both UNet and FNO predictions. As the stacking method in its most basic form can approximate soft voting by learning equal weights for each channel, it should be noted that its true potential lies in extending it with added contextual windows and nonlinearities. This allows the network to learn a nonlinear combination of the sub models’ outputs and to learn from weaknesses and strengths of them across different regions of the field. Though many more advanced ensembling techniques exist [115, 116, 117] and are considered “state of the art,” we consider them beyond the scope of this paper where our goal is to provide a set of baseline models.In Table 3, we summarize the characteristics of the three baseline models and their training schemes.
Table 3: Characteristics of the baseline models and training schemes used for comparison.
Model Parameters Input Output Activation Training scheme Loss Reference
DRM PINN 51,109 Unstructured triangular mesh $\phi,u,v$ Adaptive ReLU Transfer learning Log residual [45]
FNO 465,557 $\phi,u,v$ $\phi,u,v$ GELU Autoregressive MSE [101]
UNet 1,926,689 $\phi$ $\phi$ ReLU One-shot Dice + Focal [109]
@@
Figure 7: Figure showing: a) The initial condition and the resulting FEA simulation under the Bi-axial tensile loading and b) The results of training the PINN model (12 separate models) with three energy decomposition methods across four different seeds. We note that the individual PINN models are converging to different crack patterns. Though some of these models show plausible results, none were able to predict the correct crack pattern.
@@ 3 Results and Discussion
Figure 8: This figure presents the predicted phase field and displacement fields for each FNO model that was trained on the dataset. This shows the results for a representative sample at the $50^{th}$ percentile of performance with a Dice score of 0.6869. We can see better performing samples in Fig. 14 The top row shows the ground truth and each subsequent row shows the results of each FNO seed. Additional examples are shown in Fig. 14
@@ 3.1 Physics Informed Neural Network Baseline
Figure 9: This figure presents: a) The mean and standard deviation of the force displacement curves across all samples in the dataset, along with the three regions (elastic, initiation, propagation) indicated and b) The rollout error metrics for the FNO models, averaged across the test set and an ensemble of five models, plotted against the simulation step number. The vertical dashed lines indicate the transition between the elastic, crack initiation, and crack propagation regimes. The top panel is a dual-axis plot for the phase field($\phi$), showing the Mean Squared Error (MSE) on the left axis and the Dice score on the right. The middle and bottom panels show the MSE for the $x$ and $y$ components of the displacement field, respectively, illustrating how predictive error remains low during the elastic phase but increases significantly once cracks initiate and propagate. The rollout MSE is computed from a single, continuous prediction of the entire simulation trajectory. To start this process, each model in the ensemble is given the initial 10 steps of the simulation from the ground truth and then predicts the remaining 90 steps autoregressively. For analysis purposes only, the resulting error is then averaged within the three distinct windows shown to illustrate how error accumulates as the simulation progresses from simple to more complex behavior. The specific MSE values for each window are provided in Table 4.
@@ 3.2 Fourier Neural Operator Baseline
In Fig. 10a, we plot a histogram of the Dice score across all 40 test samples evaluated with the trained FNO models. As indicated by the vertical line labeled “Dice Mean Seed 1-5” in the Fig. 10a histogram, the average Dice score across the test set and across all seeds is 0.6208. To further investigate the performance of autoregressive models on this dataset, we divided the simulation trajectory into 3 windows, as shown in Fig 9-a. The first 20 steps corresponds to the elastic regime with no crack initiation. The second window (steps 21-40) captures the crack initiation, while the final window (steps 41-100) represents crack propagation within the domain. In Fig 9-b, we can see that the MSE of all three fields ($\phi,u_{x},u_{y}$) remains low during the elastic regime, but it increases significantly during the crack initiation phase and continues to rise during crack propagation. This is expected, as there is zero to no change in the phase field and very small variance in the displacement field during the elastic regime, making it easier for the model to approximate the solution with low error. However, as the cracks start to initiate and propagate, the phase field and the displacement field undergo significant changes, making it more challenging for the model to predict the crack evolution accurately. Figures 15-17 in the appendix provide further insight by illustrating the ground truth at the 100th step alongside the predicted crack patterns from different FNO seeds in each row and the absolute difference between these two for each channel. These results highlight both the sensitivity of FNOs to initialization, and their difficulty in capturing high frequency features such as the edges of the crack patterns. We can see in Fig. 8 that FNO models predicted a visually less well-defined and smoother phase field solution when compared to the FEA Ground Truth. A plausible explanation for this outcome is related to a well-known limitation of neural networks referred to as spectral bias, which indicates the preference of neural networks to learn low frequency features at the expense of high frequency features [26, 118, 119, 27]. Another interesting observation from Fig. 8 is that although the predicted displacement fields may deviate from the ground truth, they remain consistent with their corresponding phase fields. This suggests that the models have captured some underlying physical relationships through training on this dataset.
Table 4: Mean Squared Error (MSE) of FNO predictions averaged over trajectory windows and averaged over seeds, reported for the elastic, initiation, and propagation windows.
Data Field Elastic Window (10:20) Initiation Window (21:40) Propagation Window (41:100)
Phase Field 7.40e-04 5.08e-03 1.23e-02
X Displacement 2.54e-03 4.24e-02 7.13e-02
Y Displacement 2.65e-03 4.00e-02 7.50e-02
Looking ahead, there are several promising directions to improve upon these baseline models. These include the exploration of various data augmentation strategies, ranging from traditional techniques to more advanced generative AI based approaches. As shown in Fig. 10b, which presents crack pattern predictions ranked from best to worst based on the soft voting method (discussed further in Section 2.4), we observe that examples with lower Dice scores typically have large quantitative and qualitative deviations from the ground truth. Thus, although the top-performing predictions (e.g., those in the upper row of Fig. 10b) show strong agreement with the FEA ground truth, the substantial discrepancies in lower-performing cases highlight significant room for improvement.